OpenStudy (anonymous):

roberto purchased airline tickets for his family of 4 the tickets cost 1250 the airline also charged a luggage fee bringing the total cost to 1400 what is the percent increase in the cost due to the luggage fee

OpenStudy (anonymous):

%C=VF-VI/VI %C=1400-1250/1250 x 100% %C=0.12 x 100% %C= 12% Therefore the cost due to the luggage fee increased by 12%